OpenStudy (anonymous):

Vectors Problem:Forces of 60 pounds and 100 pounds have a resultant force of 125 pounds. Find the angle between the two forces.

OpenStudy (anonymous):

to my knowledge, one more data is missing. But still I can be wrong

OpenStudy (anonymous):

nope im sure this is all, i managed to solve this before but forgot how to

OpenStudy (anonymous):

ahaa you can represent the forces by the sides of a triangle and then it's a matter of using sine rule and cosine rule

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

so the sides are 60,100,125

OpenStudy (anonymous):

just realized law of cosines

OpenStudy (anonymous):

and the cosine rule is a^2=b^2+c^2-(bc)costheta

OpenStudy (anonymous):

yeah, so I guess you got it now

OpenStudy (anonymous):

thanks for the question as well

OpenStudy (anonymous):

ive been trying to use vector properties to solve this

OpenStudy (anonymous):

wait 1 sec the resultant force isnt the 3rd side of the triangle

OpenStudy (whpalmer4):

 \[c^2=a^2+b^2-2ab\cos\theta\]\[125^2=100^2+60^2-2(100)(60)\cos\theta\]\[2025=-12000\cos\theta\]You can do the rest...

OpenStudy (whpalmer4):

@d3v13 one of us has a faulty formulation of the law of cosines